Use the point-slope form to write an equation of the line with the given slope and point. Then write the equation in slope-intercept form.
Slope -3/2 x y
8 -3
the 8 goes under the x and the -3 goes under the y

Answer by TimothyLamb(4379)

point-slope form:
y - -3 = -3/2(x - 8)
y + 3 = -3/2(x - 8)
---
slope-intercept form:
y + 3 = -3/2(x - 8)
y + 3 = (-3/2)x + 12
y = (-3/2)x + 9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php